What is the final temperature ( 𝑇f ) of the two objects after thermal equilibrium is reached?

829 g copper pan is removed from the stove at a temperature of 150 °C. It is then stacked on top of a 611 g aluminum sheet at room temperature, 25 °C. The specific heat capacity for copper is 0.387 Jg·K and that for aluminum is 0.904 Jg·K. Assume no heat is gained or lost to the surroundings.

What is the final temperature ( 𝑇f ) of the two objects after thermal equilibrium is reached?
Answer #1

SOLUTION :


Aș per energy conservation, heat lost by copper pan is equal to heat gained by 

The aluminium sheet.


Let final temperature of pan and sheet be T º K .


So,


Heat lost by copper pan = Heat gained by the aluminium sheet

=> m coper*s copper*Temperature loss = m aluminium*s aluminium*Temp. gain.


=> 829 g * 0.387 J/(g.K) * ((150+273) - T) = 611 * 0.904 J/(g.K) * (T - (25+273))

=> 829 * 0.387 * (423 - T)  Joules = 611 * 0.904 * (T - 298)

=> 320.823 * (423 - T) = 552.344 * (T - 298)

=> (552.344 + 320.823) T = 320.823 * 423 + 552.344 * 298

=> 873.167 T = 300306.641

=> T = 300306.641 / 873.167

=> T = 343.93 ºK

=> T = 343 - 93 - 273 = 70.93 ºC 


So, final temperature will be 70.93 ºC (ANSWER).

answered by: Tulsiram Garg